A bag contains a number of 10paise,20 paise and 25 paise coins inratio 7:3:4. If value of money comes to Rs.90 then the number of 25paise coins in the bag is?
1
Expert's answer
2012-08-10T09:31:11-0400
Let the number of 25paise coins in the bag is 4*X.
Then the value of money: V
= 7*X * 0.1 + 3*X * 0.2 + 4*X * 0.25
V = 0.7*X + 0.6*X + X
V = 2.3*X =
90
4*X = approx. 157 coins